OpenStudy (anonymous):

A CD spins at a constant angular velocity of 4.0 revolutions per second clockwise. Which of the statements about the CD is true? A. No net torque acts on it at all. B. A net torque acts on it clockwise to keep it moving. C. A net torque acts on it counterclockwise to keep it moving. **Not sure! Thank you:)

OpenStudy (ipwnbunnies):

If a torque is applied, the movement should be in the same direction. If there is a rotation, or some circular motion, a torque has to be applied, I guesss. So, yeah

OpenStudy (ipwnbunnies):

movement should be in the same direction as the torque*

OpenStudy (anonymous):

ohh so the net torque would be acting on it clockwise?

OpenStudy (kainui):

We need to relate torque to velocity somehow, so let's use definitions since they are all we have: \[\Large \bar \tau = \bar r \times \bar F\] We know force is mass times acceleration, so we plug that in: \[\Large \bar F = m \bar a\] and what is acceleration? It's a change of velocity and here it says constant velocity doesn't it?

OpenStudy (ipwnbunnies):

I would believe so. I feel unsure for some reason, lol.

OpenStudy (anonymous):

yes:) @Kainui

OpenStudy (kainui):

Another way you should think of stuff is when they say "constant velocity" or constant anything is to write down that its derivative is 0\[\Large \frac{d \omega}{dt}=0\] If you don't know what derivatives are, don't worry, just think about whatever value is the changing version of that. So if they say position is constant, velocity is 0. If velocity is constant, acceleration is 0. So alternatively just as good write: \[\Large \alpha = 0\] and now you just have to relate torque to angular acceleration. =P

OpenStudy (anonymous):

ohh okay:)

OpenStudy (anonymous):

wait so it would be moving clockwise with torque based on those explanations?

OpenStudy (kainui):

Nooooo

OpenStudy (anonymous):

yes that's what i mean :)

OpenStudy (anonymous):

wait @Kainui it's not?

OpenStudy (kainui):

Since torque depends on force and force depends on acceleration, since there is 0 acceleration (constant velocity) then that means there must also be 0 torque! \[\Large \tau = r \times F = r \times (m a) \]

OpenStudy (anonymous):

oh so it's no torque is used? not torque moving clockwise?

OpenStudy (kainui):

Well I should probably be more specific in that the net torque is 0. It's the only way you can move at a constant velocity. In reality there will be a torque applied to spin it, but there will be resistance to torque from friction. But since you have achieved a constant velocity, then they have cancelled each other out.

OpenStudy (anonymous):

ahh okay:) so no net torque!

OpenStudy (ipwnbunnies):

But it's always accelerating since it's circular motion, right?

OpenStudy (anonymous):

yes circular motion! wait so it would keep moving?

OpenStudy (ipwnbunnies):

Oh, not angular velocity. I got you now, I think.

OpenStudy (kainui):

There are three types of acceleration that should try to clear up in your mind, and those are tangential, centripetal, and angular.

OpenStudy (anonymous):

ahh okay:) and this is centripetal right?

OpenStudy (anonymous):

I thought torque in a real system could never be 0

OpenStudy (kainui):

Technically we're talking about all of them right now haha. But more specifically since the problem has given you angular velocity we are "looking" at acceleration as being angular. That would be your best interpretation here. But they are all really one.

OpenStudy (anonymous):

ohh okay:) thank you guys:)

OpenStudy (anonymous):

Wait wait, no kai is right, since the `angular velocity` is `constant`, then the net torque is 0.

OpenStudy (anonymous):

haha ahh okie:) thanks for explaining all this!! :D

OpenStudy (kainui):

Well a more direct formula is: \[\Large \tau = m r^2 \alpha\]

OpenStudy (anonymous):

ohh okie:)

OpenStudy (kainui):

Since we know alpha = 0, then tau = 0.

OpenStudy (anonymous):

Bingo

OpenStudy (anonymous):

haha yes bingo:)

OpenStudy (kainui):

But of course you shouldn't ever really memorize things like that unless it's by pure accident. Remember a few formulas and then chain them together when you need to to get the results you need. Although even this is a questionable approach. In your mind you should try to imagine these scenarios and imagine what a torque really is and be able to clearly less that there can't possibly be a net torque on a constant velocity object since a torque is really a change in angular momentum. Since the mass and velocity aren't changing, there can be no torque!

OpenStudy (anonymous):

ahh okay will do! makes sense to me now, thank you so much!! :)